OpenStudy (anonymous):

How do you find the slope of the tangent lines to the given curves at the given x value, to three decimal places. a. f(x)=(1/3)^x+2 at x=0

OpenStudy (anonymous):

You have to use the rules of ln(). y = 3^x , ln(y) = ln(3^x), ln(y) = xln(3)